# Constructing and Masking Preference Profile with LLMs for Filtering Discomforting Recommendation

### v1

@@ Abstract.
discomforting recommendation filtering, recommender systems, personalized algorithm, large language model, preference profile
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Personalization
@@ 1. Introduction
In this paper, we aim to design a tool that helps users filter out discomforting recommendations. This issue has two key characteristics: (1) users’ perceptions of discomfort are highly subjective, and (2) the algorithms recommending such content operate as black-box systems. As illustrated in Figure 1, we formalize the problem as follows: black-box personalized algorithms recommend items to a user based on the inferred preference profile (often implicit in the embeddings), and our objective is to identify and filter out subjectively discomforting items for the user. Due to the opacity of algorithmic systems, the preference profile is both incomprehensible and uneditable, making it challenging for the user to influence the algorithm’s decisions.
Figure 1. Problem formalization.
[…]
@@
Given the natural language understanding, reasoning, and generation capabilities demonstrated by Large Language Models (LLMs) (Li et al., 2024; Zhao et al., 2023), we propose that LLM provide a promising solution for achieving these design goals. To this end, we designed an LLM-based tool named DiscomfortFilter, specifically aimed at helping users filter out discomforting recommendations (Section 4). Figure 2 illustrates the workflow of DiscomfortFilter: (1) DiscomfortFilter observes algorithm-recommended items to construct a comprehensible and editable preference profile for the user; (2) Through a guided conversation, DiscomfortFilter assists the user in expressing personalized filtering needs, and then masks the discomforting preferences with in the profile; (3) DiscomfortFilter filters out discomforting recommendations based on the edited preference profile in a plug-and-play manner, ensuring that user-perceived items no longer includes discomforting elements. Additionally, DiscomfortFilter provides the user with access to filtering logs recorded during step (3), assisting the user in refining filtering needs in a manner similar to step (2). Overall, DiscomfortFilter empowers the user to actively influence the decisions made by personalized algorithms, enhancing control over the algorithms.
@@
We validated the efficacy of the constructed preference profile using an offline proxy task. The findings suggest that it enhances the reasoning process of LLMs, markedly decreases the challenge of aligning them with users, and allows a 3.8B open-source LLM to rival top commercial models. Additionally, we conducted a one-week user study on Zhihu (a platform similar to Quora), China’s largest Q&A community, with 24 participants (Section 5). The results demonstrate that our design goals effectively help users express their filtering needs and filter out discomforting recommendations, with DiscomfortFilter successfully achieving these goals. We also analyzed how DiscomfortFilter impacts platform recommendation outcomes by influencing the exposure of discomforting items.
Finally, we conducted an in-depth discussion (Section 6), covering: challenges of filtering discomforting recommendations with LLMs, relevance to research topics in recommender systems, potential impact on platforms, and limitations and future work.
The key contributions of this study are outlined below:
@@
  • We conducted a formative study to examine the current status and user expectations regarding the filtering of discomforting recommendations.
@@
  • We designed and introduced a LLM-based tool called DiscomfortFilter to assist users in filtering out discomforting recommendations. To the best of our knowledge, our study represents the initial investigation into the application of LLMs for this specific task.
  • We demonstrated the efficacy of DiscomfortFilter through an offline proxy experiment and a user study.
  • We engaged in an in-depth discussion of this study, thereby enhancing its broader implications.
@@ 3.1. Findings from Semi-Structured Interviews
F1: Users may encounter discomforting recommendations for three reasons. (1) User behavior deviation. Curiosity-driven search behavior and clickbait-induced clicks may fail to reflect a user’s true long-term interests, leading inaccurate user preference modeling. For example, P03 said “Out of curiosity, I once searched for adult products, and now they keep showing up in my recommendations—so embarrassing.(2) Algorithmic modeling bias. Personalized algorithms cannot fully capture the nuanced interests [^2] This stems from collaborative filtering only retaining low-frequency information (Shen et al., 2021). and contexts of users. For example, P06 said “Getting horror content at night is awful, even if I watch it during the day.(3) Conflicting interests. For instance, platforms may promote content designed to boost user engagement, even if it may cause discomfort. Seven participants mentioned scenarios where this was the case.
@@ 4.1. Overview
The upper part of Figure  4 illustrates how personalized algorithms present items to the user after the introduction of DiscomfortFilter. By integrating a Content Filter Module into the original items presentation flow, DiscomfortFilter removes discomforting recommendations, ensuring that only non-discomforting items are ultimately displayed to the user. The identification of discomfort is based on user-configured filtering rules, giving the user a sense of control over the process.
@@ 4.2.2. Preference Profile Construction Module
As illustrated in Figure 5, we propose an LLM-based multi-agent pipeline to complete this process. We adopt the general assumption of pairwise rank learning—when two items are displayed simultaneously, the one that is clicked is more appealing to the user. For each clicked item (denoted as pos for positive), the module randomly samples an unclicked item that appeared simultaneously with pos as a negative item (denoted as neg), and then constructs an ordered pair ¡pos, neg¿. Note that pos and neg are unprocessed raw contents. Then, this pair is processed by the pipeline.
Figure 5. The User Profile Construction Module is a multi-agent pipeline powered by LLMs.
[…]
First, the Perceive Agent identifies pos and neg from the user’s perspective, analyzing why the user clicked on pos but not on neg based on the current preference profile. Different users focus on different aspects of the same item. By incorporating a user’s preference profile, the Perceive Agent can accurately identify the aspects that matter most to each individual.
@@
Finally, the feature nodes are ranked using the PageRank algorithm. Features with higher rankings are generally more aligned with the user’s interests, while lower-ranked features tend to be less relevant.
@@ 5.1.2. Dataset
We conducted offline experiments using the MIND (Wu et al., 2020) dataset, which details negative samples recommended to users during interactions, along with side information. Given the cost of API calls, we sampled a subset of users for our experiment. In order to account for the varying interaction frequencies, we first grouped users by interaction frequency into intervals: $[0,10)$, $[10,20)$, and up to $[100,+\infty)$. Then, from each group, we randomly selected users, ensuring a total of at least 10,000 interactions per group.
Table 1. Confusion matrix.
Predicted Positive Predicted Negative
Actual Positive 173 32
Actual Negative 67 208
@@ 5.1.4. Results
The profiles constructed by several ablation baselines require sufficiently powerful LLMs to achieve better performance, leading commercial models to outperform open-source LLMs on these baseline methods. However, our construction method reduces task complexity by streamlining the reasoning process through a well-constructed preference profile. This enhancement enables open-source models to match or even surpass the performance of certain commercial models.
Figure 7. Score distribution from the questionnaire: the integers in the colored bars represent the number of participants for each score, with the color-to-score relationship indicated in the bottom-right corner
[…]
@@ 5.2.4. Results of the questionnaire
Guided by the Technology Acceptance Model (TAM), we developed three evaluation questions for each design goal and the overall tool, focusing on perceived usefulness (PU), perceived ease of use (PEOU), and behavioral intention (BI). Figure 7 shows the specific questions and the corresponding score distributions. For each question, at least three-quarters of participants rated it 4 or 5, indicating overall satisfaction with the role of DiscomfortFilter in assisting users to filter out discomforting recommendations. With a Cronbach’s $\alpha$ of 0.80, we believe the results are reliable for subsequent analysis.
@@ 5.2.5. Results of the interview
R1. Natural language filtering rules helped participants in personalizing their filtering needs (G1). Some participants also found emotional value in being able to “complain to the tool about the platform’s recommendations” (P31) and “set rules regarding mood” (P38). However, false associations in LLMs sometimes hindered accurate identification of discomforting recommendations. Most participants noted that DiscomfortFilter occasionally overextended the rules, leading to unintended content filtering. This reduced precision — as shown in Table 1, where 24 participants annotated 480 recommended contents, resulting in a precision rate of 173/240=72.1% and a recall rate of 173/(173+32)=84.4%.
R2. Providing preference explanations helps participants identify and articulate their filtering needs (G2). Participants were sometimes unclear about their filtering needs, and the preference explanations enabled them to “discuss with the tool how to establish rules” (P31). However, some participants felt that the preference profile lacked sufficient detail, offering only a “broad overview of the recommended content” (P17).
@@
R5. The plug-and-play approach facilitates flexible configuration of dynamic filtering needs (G4), adapting to contextual changes and short-term interests. However, some participants expressed dissatisfaction with the absence of an efficient method for managing numerous filtering rules, noting that “reviewing all filtering rules to decide which to activate is cumbersome” (P33).
@@ 5.2.7. Impact on platform recommendation outcomes
Assuming DiscomfortFilter processes $N$ items using a filtering rule, with $n$ items identified as discomforting, we define $n/N$ as the filtering efficiency of this rule. We calculated the daily average filtering efficiency for all filtering rules that remained active for more than five days during the seven-day user study. The trend of daily average filtering efficiency from the day they were configured is shown in Figure 8. Over time, the filtering efficiency steadily declined, indicating that the platform was recommending progressively fewer discomforting items. This decline can be attributed to the introduction of DiscomfortFilter, which reduces users’ exposure to discomforting recommendations, thereby resulting in fewer interactions with such items over time. As a result, the platform’s recommender systems can dynamically adjust to users’ evolving preferences, further reducing the likelihood of discomforting items being recommended. It is important to emphasize that, from the users’ perspective, discomforting recommendations vanish immediately, as they were filtered out entirely.
Figure 8. Filtering Efficiency Over 5 Days.
[…]
@@ 6.4. Limitations and Future Work
7. Conclusions
Building on insights from a formative study, we developed an LLM-based tool called DiscomfortFilter to assist users in identifying and filtering discomforting recommendations. Results from an offline experiment and a user study demonstrated DiscomfortFilter’s effectiveness. In-depth discussions illuminated future directions and the potential broad impact of our work. We believe our study can strengthen the recommender systems community’s focus on human-centered design.
@@ C.2.1. Perceive Agent
Now, for a question titled ¡pos¿/¡neg¿, assuming you have/have not interacted with it, please explain your reasons.
@@ C.2.2. Summary Agent
A Zhihu user has/has not interacted with the question titled ¡title¿, and here are the reasons she/he provided: ¡reasons¿. Based on the reasons, please summarize the key features of the question.
@@ E.3. LLMs as Personal Assistants
Several commercial products, such as Microsoft’s Copilot, New Bing, Google’s Bard, and Gemini, as well as offerings from smartphone manufacturers, have integrated LLMs to boost productivity and enhance web and on-device experiences (Mehdi, 2023a; Achiam et al., 2023; Mehdi, 2023b). In the area of human-centered recommender systems, the RAH framework (Shu et al., 2024), a LLM-based multi-agent system, analyzes user preferences by observing ratings and reviews and incorporates a reflection mechanism to improve the accuracy. Recent research has explored the use of LLMs as personal assistants for various special populations. Glasser et al. (2022) examined how deaf and hard-of-hearing users interact with LLM-based personal assistants capable of understanding sign language to better customize these technologies for their needs. Additionally, some work focuses on designing interfaces for visually impaired individuals to access visual information using LLMs and visual language models (Zhang, 2024). Finally, Jang et al. (2024) investigated how autistic workers use LLMs for communication assistance, highlighting their reliance on and trust in these tools.

### v2

# Filtering Discomforting Recommendations with Large Language Models

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In this paper, we aim to design a tool that helps users filter out discomforting recommendations. This task has two key challenges: (1) users’ perceptions of discomfort are highly subjective, and (2) the algorithms recommending such content operate as black-box systems. As illustrated in Figure 1, we formalize the problem as follows: black-box personalized algorithms recommend items to a user based on the inferred preference profile (often implicit in the embeddings), and our objective is to identify and filter out subjectively discomforting items for the user. Due to the opacity of algorithmic systems, the preference profile is both incomprehensible and uneditable, making it challenging for the user to influence the algorithm’s decisions.
Figure 1. Problem formulation.
Figure 2. The workflow of DiscomfortFilter.
@@
Given the natural language understanding, reasoning, and generation capabilities demonstrated by LLMs (Li et al., 2024; Zhao et al., 2023), we propose that LLM provide a promising solution for achieving these design goals. To this end, we designed an LLM-based tool named DiscomfortFilter, specifically aimed at helping users filter out discomforting recommendations (Section 4). Figure 2 illustrates the workflow of DiscomfortFilter: (1) DiscomfortFilter identifies algorithm-recommended items based on the user’s personalized perceptions, integrates the user’s pairwise preferences, and ranks them to construct a comprehensible and editable preference profile tailored for the user; (2) Through a guided conversation, DiscomfortFilter assists the user in expressing personalized filtering needs, and then masks the discomforting preferences with in the profile; (3) DiscomfortFilter filters out discomforting recommendations based on the edited preference profile in a plug-and-play manner, ensuring that user-perceived items no longer includes discomforting elements. Additionally, DiscomfortFilter provides the user with access to filtering logs recorded during step (3), assisting the user in refining filtering needs in a manner similar to step (2). Overall, DiscomfortFilter empowers the user to actively influence the decisions made by personalized algorithms, enhancing control over the algorithms.
@@
We validated the efficacy of the constructed preference profile using an offline proxy task. The findings suggest that it enhances the reasoning process of LLMs, markedly decreases the challenge of aligning them with users, and allows a 3.8B open-source LLM to rival top commercial models. Additionally, we conducted a one-week user study on Zhihu (a platform similar to Quora), China’s largest Q&A community, with 24 participants (Section 5). The results demonstrate that our design goals effectively help users express their filtering needs and filter out discomforting recommendations, with DiscomfortFilter successfully achieving these goals. We also analyzed how DiscomfortFilter impacts platform recommendation outcomes by influencing the exposure of discomforting items. Finally, we conducted an in-depth discussion (Section 6), covering: challenges of filtering discomforting recommendations with LLMs, relevance to research topics in recommender systems, potential impact on platforms, and limitations and future work.
The key contributions of this work are outlined below:
@@
  • We conducted a formative study with 15 participants to examine the current status and user expectations regarding the filtering of discomforting recommendations.
@@
  • We designed an LLM-based tool named DiscomfortFilter to assist users in filtering out discomforting recommendations, which is the first attempt to leverage LLMs in this important task, to the best of our knowledge.
  • We evaluated DiscomfortFilter through an offline proxy experiment and a user study and the results showed that DiscomfortFilter can effectively help users express their filtering needs and filter out discomforting recommendations.
  • We discussed the challenges and opportunities of using LLMs for filtering discomforting recommendations and shed light on its broader implications.
@@ 3.1. Findings from Semi-Structured Interviews
F1: Users may encounter discomforting recommendations for three reasons. (1) User behavior deviation. Curiosity-driven search behavior and clickbait-induced clicks may fail to reflect a user’s true long-term interests, leading inaccurate user preference modeling. For example, P03 said “Out of curiosity, I once searched for adult products, and now they keep showing up in my recommendations—so embarrassing.(2) Algorithmic modeling bias. Personalized algorithms cannot fully capture the nuanced interests [^2] This stems from the fact that collaborative filtering algorithms mainly retain low-frequency information during model training (Shen et al., 2021). and contexts of users. For example, P06 said “Getting horror content at night is awful, even if I watch it during the day.(3) Conflicting interests. For instance, platforms may promote content designed to boost user engagement, even if it may cause discomfort. Seven participants mentioned scenarios where this was the case.
@@ 4.1. Overview
The upper part of Figure  4 illustrates how personalized algorithms present items to the user after the introduction of DiscomfortFilter. By integrating a Content Filter Module into the original items presentation flow, DiscomfortFilter removes discomforting recommendations, ensuring that only non-discomforting items are ultimately displayed to the user. The identification of discomfort is based on user-configured filtering rules, giving the user an ability of control over the process.
@@ 4.2.2. Preference Profile Construction Module
As illustrated in Figure 5, we propose an LLM-based multi-agent pipeline to complete this process. We adopt the general assumption of pairwise rank learning (Rendle et al., 2012)—when two items are displayed simultaneously, the one that is clicked is more appealing to the user. For each clicked item (denoted as pos for positive), the module randomly samples an unclicked item that appeared simultaneously with pos as a negative item (denoted as neg), and then constructs an ordered pair ¡pos, neg¿. Note that pos and neg are unprocessed raw contents. Then, this pair is processed by the pipeline.
Figure 5. The Preference Profile Construction Module is a multi-agent pipeline powered by LLMs.
First, the Perceive Agent identifies pos and neg from the user’s perspective, analyzing why the user clicked on pos but not on neg based on the current preference profile. Different users focus on different aspects of the same item. By incorporating a user’s preference profile, the Perceive Agent can accurately identify the aspects that matter most to each individual user.
@@
Finally, the feature nodes are ranked using the PageRank algorithm, which provides a comprehensive ranking of preferences. Features with higher rankings are generally more aligned with the user’s interests, while lower-ranked features tend to be less relevant.
@@ 5.1.2. Dataset
We conducted offline experiments using the MIND dataset (Wu et al., 2020), which details negative samples recommended to users during interactions, along with side information. Given the high cost of API calls, we sampled a subset of $\sim$5,000 users for experimentation. To account for the varying interaction frequencies, we first grouped users by interaction frequency into intervals: $[0,10)$, $[10,20)$, and up to $[100,+\infty)$. Then, from each group, we randomly selected users, ensuring a total of at least 10,000 interactions per group.
@@ 5.1.4. Results
The profiles constructed by several ablation baselines require sufficiently powerful LLMs to achieve better performance, leading commercial models to outperform open-source LLMs on these baseline methods. However, our construction method reduces task complexity by streamlining the reasoning process through a well-constructed preference profile. This enhancement enables open-source models, e.g., Phi-3, to match or even surpass the performance of certain commercial models, e.g., GPT-4o-mini.
Figure 7. Score distribution from the questionnaire: the integers in the colored bars represent the number of participants for each score, with the color-to-score relationship indicated in the bottom-right corner.
@@ 5.2.4. Results of the questionnaire
Guided by the Technology Acceptance Model (TAM) (Davis et al., 1989; Marangunić and Granić, 2015), we developed three evaluation questions for each design goal and the overall tool, focusing on perceived usefulness (PU), perceived ease of use (PEOU), and behavioral intention (BI). Figure 7 shows the specific questions and the corresponding score distributions. For each question, at least three-quarters of participants rated it 4 or 5, indicating overall satisfaction with the role of DiscomfortFilter in assisting users to filter out discomforting recommendations. With a Cronbach’s $\alpha$ of 0.80, we believe the results are reliable for subsequent analysis.
@@ 5.2.5. Results of the interview
R1. Natural language filtering rules helped participants in personalizing their filtering needs (G1). Some participants also found emotional value in being able to “complain to the tool about the platform’s recommendations” (P31) and “set rules regarding mood” (P38). However, false associations in LLMs sometimes hindered accurate identification of discomforting recommendations. Seven participants noted that DiscomfortFilter occasionally overextended the rules, leading to unintended content filtering. This reduced precision — as shown in Table 4, where 24 participants annotated 480 recommended contents, resulting in a precision rate of 173/240=72.1% and a recall rate of 173/(173+32)=84.4%.
R2. Providing preference explanations helps participants identify and articulate their filtering needs (G2). Participants were sometimes unclear about their filtering needs, and the preference explanations enabled them to “discuss with the tool how to establish rules” (P31). However, two participants felt that the preference profile lacked sufficient detail, offering only a “broad overview of the recommended content” (P17).
@@
R5. The plug-and-play approach facilitates flexible configuration of dynamic filtering needs (G4), adapting to contextual changes and short-term interests. However, three participants expressed dissatisfaction with the absence of an efficient method for managing numerous filtering rules, noting that “reviewing all filtering rules to decide which to activate is cumbersome” (P33).
@@ 5.2.7. Impact on platform recommendation outcomes
Assuming DiscomfortFilter processes $N$ items using a filtering rule, with $n$ items identified as discomforting, we define $n/N$ as the filtering burden of this rule. We calculated the daily average filtering burden for all filtering rules that remained active for more than five days during the seven-day user study. The trend of daily average filtering burden from the day they were configured is shown in Figure 8. Over time, the filtering burden steadily declined, indicating that the platform was recommending progressively fewer discomforting items. This decline can be attributed to the introduction of DiscomfortFilter, which reduces users’ exposure to discomforting recommendations, thereby resulting in fewer interactions with such items over time. As a result, the platform’s recommender systems can dynamically adjust to users’ evolving preferences, further reducing the likelihood of discomforting items being recommended. It is important to emphasize that, from the users’ perspective, discomforting recommendations vanish immediately, as they were filtered out entirely.
Figure 8. Average filtering burden over 5 days.
@@ 6.4. Limitations and Future Work
7. Conclusion
Building on insights from a formative study, we developed an LLM-based tool named DiscomfortFilter to assist users in identifying and filtering discomforting recommendations from recommender systems. Results from an offline experiment and a user study demonstrated DiscomfortFilter’s effectiveness. In-depth discussions illuminated future directions and the potential broad impact of our work. We believe our study can strengthen the recommender systems community’s focus on human-centered design.
@@ Acknowledgements.
Sincere thanks to all participants involved in this work.
@@ C.2.1. Perceive Agent
Now, for a question titled ¡title¿, assuming you have (or have not) interacted with it, please explain your reasons.
@@ C.2.2. Summary Agent
A Zhihu user has (or has not) interacted with the question titled ¡title¿, and here are the reasons she/he provided: ¡reasons¿. Based on the reasons, please summarize the key features of the question.
@@ E.1. Discomforting Recommendations
Table 4. Confusion matrix.
Predicted Positive Predicted Negative
Actual Positive 173 32
Actual Negative 67 208
@@ E.3. LLMs as Personal Assistants
Several commercial products, such as Microsoft’s Copilot, New Bing, Google’s Bard, and Gemini, as well as offerings from smartphone manufacturers, have integrated LLMs to boost productivity and enhance web and on-device experiences (Mehdi, 2023a; Achiam et al., 2023; Mehdi, 2023b). In the area of human-centered recommender systems, the RAH framework (Shu et al., 2024), an LLM-based multi-agent system, analyzes user preferences by observing ratings and reviews and incorporates a reflection mechanism to improve the accuracy. Recent research has explored the use of LLMs as personal assistants for various special populations. Glasser et al. (2022) examined how deaf and hard-of-hearing users interact with LLM-based personal assistants capable of understanding sign language to better customize these technologies for their needs. Additionally, some work focuses on designing interfaces for visually impaired individuals to access visual information using LLMs and visual language models (Zhang, 2024). Finally, Jang et al. (2024) investigated how autistic workers use LLMs for communication assistance, highlighting their reliance on and trust in these tools.
Appendix F More Discussion
Here are some discussions that help readers gain a deeper understanding of our work.
F.0.1. Generalizability to non-text-based items
In Section 6.2, we suggested future extensions to multimodal content. Expanding to non-text scenarios is a natural progression, requiring the use of multimodal models to understand such content. However, the efficiency of processing, especially for videos, remains a concern. Preprocessing non-text content offline may offer a viable solution.
F.0.2. Differences between discomforting and disliked items
We believe this primarily manifests in the impact perceived by users, with “discomforting” items having a greater impact. For uninteresting or disliked items, users may simply choose to ignore them. However, as we mentioned in the introduction, discomforting recommendations may not only fail to engage users but also lead to negative emotional consequences, such as anxiety, unease, or distress. This is a highly subjective matter and has a dynamic nature. For example, as illustrated in our formative study (Section 3.1, F1), one participant mentioned that he might search for horror content to watch during the day, but encountering the same horror content at night would make him feel very uncomfortable, even to the extent of affecting his sleep.
In our formative study, we observed that most users do not express strong hostility toward uninteresting (disliked) content but are more averse to “discomforting” content. Thus, from the users’ perspective, filtering discomforting content is a more pressing issue than merely addressing disinterest, and this study focuses on addressing the former.
F.0.3. Discussion of critiquing-based recommender systems.
Critiquing-based systems (Wu et al., 2019; Toroghi et al., 2023) adjust recommendations end-to-end based on feedback, whereas our work adopts a plug-and-play filtering approach, enabling immediate real-world applicability. As noted in Section 6.2, our work complements algorithmic designs like critiquing systems.
F.0.4. Computational overhead and scalability.
There are three areas where LLMs are needed: (1) to determine whether an item should be filtered, (2) for conversational agents, and (3) to build preference profiles. Among these, (1) and (3) have higher real-time requirements.
For (1), assume a user has $m$ filtering rules. Processing each item requires $1+m$ LLM calls, including one call for analyzing the item and $m$ calls to check individually whether the item should be filtered according to each filtering rule. In the user study, we recruited 24 participants, and they did not notice any significant delay with the LLM service during their use.
For (3), suppose the algorithm recommends $N$ items, and a user clicks on $n$ of them ($N\gg n$). One direct way to construct a user’s preference profile is to analyze all the items recommended by the algorithm, treating the clicked items as positive samples and the unclicked ones as negative samples. This process is similar to pointwise matrix factorization in collaborative filtering, with a complexity of $O(N)$. However, we adopted the pairwise ranking approach, where for each positive sample clicked by the user, a negative sample is sampled, reducing the complexity to $O(2n)$ (with $n$ positive samples and corresponding $n$ negative samples). This process is similar to the matrix factorization of BPR Loss in collaborative filtering with a negative sampling ratio equal to 1. Additionally, in the user study, we found that users could smoothly use the tool without experiencing delays, indicating good real-time performance.
We are working hard to publish our work to the extension store of the Chrome browser, where users can access commercial LLM services through a private API Key. Meanwhile, we are exploring collaborations with commercial partners, and when our work is applied to real-world applications, certain engineering designs will be considered, such as processing items offline and using different sizes of LLMs for different tasks. In the future, we believe that all our services, including LLM services, have the potential to run entirely on mobile devices. This is also why we chose to conduct offline experiments with the Phi series LLM, which has 3.8 billion parameters. From its inception, the Phi series model was designed with running on mobile devices in mind, and we believe that LLMs on user devices will become a reality in the near future.
F.0.5. Discussion on the echo chamber problem.
Research indicates that merely providing users with diverse recommendations is not sufficient to avoid the echo chamber problem (Zhang et al., 2024). Active exploration by users is key to addressing the echo chamber effects (Zhang et al., 2024). We believe that, through actively guiding users, the critiquing-based recommender systems, as well as our work, can contribute to achieving this goal.
Our work provides users with understandable and editable preference profiles, giving them greater agency over the recommended content. In this context, if systems similar to the critiquing-based recommender systems can effectively capture users’ preferences in real time and then deliver diverse recommendations, it becomes possible to prevent the formation of information silos. This discussion further underscores why we believe our human-centered approach complements algorithm-centered work in the research and development of recommender systems.
F.0.6. It would be helpful to compare the effectiveness of the filtering approach presented in this paper with those used in other studies on recommendation unlearning.
Recommendation unlearning emphasizes the model’s capability, primarily focusing on the completeness, efficiency, and performance of the model after forgetting. These methods typically predefine which interactions should be forgotten, assess the completeness of forgetting through Membership Inference Attacks (MIA), measure the efficiency of forgetting through runtime analysis, and evaluate post-forgetting performance through recommendation accuracy. In contrast, DiscomfortFilter filters interactions from the perspective of user needs—due to its subjectivity, it is difficult to clearly define which interactions should be filtered in advance. Therefore, it is hard to make a quantitative comparison with recommendation unlearning, as current unlearning methods cannot handle subjective user inputs. We believe that these two types of work focus on model design and human-centered design, respectively, and are complementary to each other.
F.0.7. It is unclear how well this effect will work for ephemeral filtering criteria.
We first selected filtering rules with more than 600 items processed. For each group of 30 items, we calculated the filtering burden of the rules. The results showed that the average filtering burden of the rules decreased in oscillations, and the corresponding fitted curve yielded conclusions similar to those from the analysis at a daily granularity.
F.0.8. Only the MIND dataset was used in the offline experiment.
As shown in Figure 5, DiscomfortFilter constructs positive and negative sample pairs by observing the content recommended by the algorithm and the user’s click behavior on that content. Specifically, our negative sampling is not performed randomly from the entire item set; instead, it is selected from items that appeared alongside the clicked content but were not clicked. This approach places high demands on the dataset and is often difficult to satisfy. Among commonly used datasets, we found that only the MIND dataset provides both unclicked items that appeared alongside the clicked ones and side information for those items. Additionally, some less common datasets, such as the ZhihuRec-Dataset, provide unclicked items but lack detailed side information, offering only token vectors. Furthermore, while classic CTR datasets include labels indicating whether a user clicked on an item, they are not suitable for our purpose, as their features are anonymous and do not contain textual information.
To address your concerns, we have relaxed the stringent requirements for the offline experiments on the dataset and included two additional datasets for validation. We consider items that are interacted with within a time window to be simultaneously recommended to a user. Items with a rating greater than or equal to 4 were considered clicked, while those with a rating below 4 were considered unclicked. We conducted experiments on the ML-1M and Amazon Book datasets, and the results were consistent with those obtained from the MIND dataset.
F.0.9. Insufficient diversity of platform and user in user study
During the user study, we ceased recruiting participants when we observed that additional participants were no longer providing new insights. Due to the substantial time investment required for user research (including interviews, transcription, coding, and analysis), it became challenging to expand the study further. For instance, two related studies recruited 15 and 12 participants, respectively (Gu et al., 2021; Smith et al., 2023). We are currently trying to collaborate with industry partners to implement the DiscomfortFilter concept in real-world recommender systems, which will allow for broader research in this area.
F.0.10. Can DiscomfortFilter be used to capture interest?
We believe the process can be understood as re-ranking algorithmically recommended content on the user side. We agree that the DiscomfortFilter could indeed be applied for such purposes and believe this might be a very promising direction for the future.
F.0.11. Is it possible to use the “Not Interested” button to enhance the feedback?
While we agree that the “Not Interested” button can enhance feedback, we find it inappropriate for our goals. In the formative study, we identified drawbacks of the “Not Interested” button, such as its lack of personalization, which prevents users from providing nuanced feedback, and its lack of flexibility, which can result in content disappearing permanently even if the disinterest is temporary. Including such a button could introduce uncontrollability and potentially harm users.
F.0.12. DiscomfortFilter or DisinterestFilter?
In our formative study, we observed that most users do not have strong hostility toward uninteresting content but are more averse to “discomforting” content. Thus, from users’ perspective, filtering discomforting content is a more pressing issue than merely addressing disinterest, and the study focuses on the former. That said, DiscomfortFilter could also be used to filter uninteresting content—the tool’s functionality depends on the user.
F.0.13. The A/B test for the control group.
Our findings revealed that DiscomfortFilter’s design inherently accounts for contestability, ensuring transparency and user control over its decisions. For instance, users can review filtering log and understand the rationale behind each filtering record, with the ability to make adjustments if the filtering records do not meet their expectations. Consequently, when we presented users with a mock filtering interface—where no actual or only random filtering occurred—they quickly discerned they were part of the control group, leading to immediate expressions of dissatisfaction. Moreover, during the user study in current version, participants expressed satisfaction with the performance of the functionalities, not just their design (Section 5.2.4 & 5.2.5). In summary, we posit that the mere presence of a filtering interface is unlikely to enhance user satisfaction. Instead, satisfaction levels are significantly influenced by the efficacy of the actual filtering performance.
Regarding other design elements, such as explanations of preference profiles and filtering rules, these strategies emerged from our participatory design process as necessities identified by participants. Furthermore, during the evaluation phase, these features garnered high satisfaction in both questionnaires and interviews with regards to their effects and design. Therefore, we believe that our current conclusions sufficiently demonstrate the effectiveness of these designs without necessitating a control group experiment.
In conclusion, the results of the A/B test do not alter our original conclusions.